# Journal Title

### v1

@@ Abstract
We use direct numerical simulation (DNS) to investigate mass transfer between liquid steel and slag during a metallurgical secondary refinement process through two reduced-scale water experiments, which reproduce the dynamics seen in an industrial bottom-blown ladle. A container is filled with water and topped by a thin layer of oil, representing the molten steel and slag, respectively. The system is agitated by a bubble plume that impinges on the oil layer and forms an open-eye. A tracer species, dissolved in the water, acts as a passive scalar that is progressively absorbed into the oil layer. Both the hydrodynamics and mass transfer in the system are studied and compared with experiments from the literature of different size and geometry.
The numerical simulation of mass transfer is challenging due to the high Péclet number, leading to extremely thin species boundary layers at the interface. Resolving the boundary layer is prohibitive even with adaptive grid techniques. A subgrid-scale (SGS) boundary layer model corrects the scalar transport equation, allowing us to solve convection-dominated transport on relatively coarse grids. The hydrodynamics is investigated, and we analyze how the resultant flow field governs mass transport. The numerical results recover two flow regimes: a quasi-steady regime at low flow rates with small deformations of the oil-water interface and an atomizing regime at large flow rates. Interfacial species transport is determined to be dominated in an annulus surrounding the open eye caused by a shear layer at the oil-water interface. It is observed that we achieve grid-independent macroscopic quantities that match relatively well with those observed in experiments, allowing use of simulation techniques as a complementary tool going forward.
@@ 1 Introduction
Performing experimental trials in the metallurgical industry is challenging because of high temperatures and the opaque nature of molten metals, consequently requiring specialized measurement techniques. For these reasons, industrial trials on the actual reactors are very expensive and practically unfeasible. Experimental campaigns are hence generally performed on reduced-scale model ladles and then a similarity criterion is used to grant a relation with the larger industrial ladles. While some studies use liquid metals at lower temperatures [3, 4, 5] some others further simplify the problem using fluids such as water and oil at room temperature [6, 7, 8]. In this case, a tracer species that is miscible in oil and water is chosen that mimics the species transfer kinetics observed in the industrial ladle. Thymol is often used as the diffusive tracer in the performed experiments.
Measuring local quantities in the fluid system requires more sophisticated measurement techniques, so the flow has historically been studied by measuring and analyzing a few global quantities: the open-eye size, the fragmentation of the slag layer and the rate of mass transfer. Numerical simulations, on the other hand, can greatly improve the understanding of the fluid dynamics in gas-stirred ladles since they allow for easier measurements of local quantities and better visualisation of the flow behaviour. While the first simulations on the topic date more than 20 years back [9, 10], with the computational power available today increasingly complex simulations have become possible [11]. In particular, while the earliest simulations mainly focused on the flow patterns in the ladle, linking it to the mass transfer by means of models and theories, more recent studies started to address the mass transfer problem itself [12, 8]. There are however still some gaps in the understanding of the problem. Furthermore, most of the studies involving numerical simulations do not resolve the turbulent flow explicitly, but they rather use turbulence models such as large eddy simulations or Reynolds-averaged Navier-Stokes to treat smaller scale phenomena. Only recently some studies on ladles have been performed using direct numerical simulations (DNS) that solve turbulence to the smallest scales [13, 14]. Lastly, the majority of numerical studies has focused on the hydrodynamic problem and ignores resolving mass transfer.
@@ 2.1 Model description
$\overline{Sh}=\frac{K_{\mathrm{w}}Ah_{\mathrm{w}}}{L_{\mathrm{x}}^{2}D_{\mathrm{w}}}.$ (2)
@@ 2.2 Mathematical model
The experiments analyzed are isothermal and they involve three immiscible phases: water, air and oil. While water and oil are incompressible, the injected air bubbles are in theory compressible. To simplify the model and reduce the computational load, we neglect the compressibility effects of air. This hypothesis is reasonable if the hydrostatic pressure of the system is significantly smaller than the atmospheric pressure $\rho_{\mathrm{w}}gh_{\mathrm{w}}\ll p_{\mathrm{atm}}$. We have to mention that, as shown in Table 1, we choose to use a value of the air density $\rho_{\mathrm{a}}$ that is ten times bigger than the air density at STP. It was observed that this change does not significantly impact the results of the simulations. However, a reduction of the density ratio $\rho_{\mathrm{w}}/\rho_{\mathrm{a}}$ makes the pressure-Poisson problem converge significantly faster, increasing computational efficiency. Interface capturing is included with the multi-VOF approach, with three indicator functions $f_{i}$ to identify the different phases. The $f_{i}$ have to respect a volume conservation equation
@@
Under the formulated hypotheses, we write the Navier-Stokes equations for the three-phase problem using the one-fluid method as
@@
Furthermore, Henry’s law holds, stating that the ratio between the oil and water-side concentration is given by a partition coefficient $P$
@@
Numerically, the scalar transport equation is solved using a two-field approach where we transport the conserved quantities
@@
corresponding to the species mass in the oil and water phases. The phasic scalar transport equation for $g$ is written as
@@
where $A_{i}$ is the phasic area. The diffusive flux across the interface is included as equal and opposite explicit source terms in the transport equation.
@@ 2.3 Simulations setup
Furthermore, the Joubert simulations underestimated the spreading force at the three-phase contact line, which caused the open-eye area to be over-predicted by up to 30 precent compared to experiments. The improved surface tension calculation due to the hybrid curvature calculation described in 2.2 led to the formation of a thin oil meniscus that significantly reduced the size of the open eye, physically consistent with the positive spreading coefficient $S_{o}=\sigma_{aw}-\sigma_{ao}-\sigma_{ow}$ of the oil-water-air system. However, meniscus coverage was observed to be grid dependent as coarse grids could not capture the thin film. Furthermore, comparisons to experiments would be difficult in the presence of thin film as sophisticated optical tools were not used when measuring the open-eye in experiments. Using the meniscus area coverage as an error bar on the open-eye area in simulations led to relative agreement with experiments (see [28]).
@@ 2.3.4 Passive Scalar Subgrid-Scale modelling
To derive a subgrid model, we first assume that the mass transfer resistance in the water layer is much larger than that of the oil layer, known from the global kinetics of the system. This allows us to compute the interfacial flux with a one-sided subgrid model in the water phase. We consider an advection-diffusion problem entirely located inside the layer $|z-z_{ow}|\ll\delta_{u,w}$, solved in the reference frame of the moving interface. In this sublayer, the velocity field $U$ may be linearized so that $U_{r}(z,t)=U_{\Sigma}(t)+\omega(t)(z-z_{ow})$, representing an interface slip velocity and a shear dependent term. We ignore tangential diffusion. The described boundary layer approximation leads to the simplified advection-diffusion equation for the concentration field $C(r,z,t)$
@@ 3 Results
This section presents the simulation results for both the hydrodynamics and mass transfer. Detailed results are provided for the Kim & Fruehan ladle, while only principal results are highlighted for the Joubert case, for which additional plots and figures are available in the supplementary materials.
@@ 3.2 Hydrodynamics results
Fig. 3 shows the conical ladle of the Kim & Fruehan case from a side view for the highest and lowest air injection rates investigated, namely 0.5 and 5.0 l/min, at three subsequent time steps. The air surface is coloured in green and the oil surface in red. The first snapshot of each simulation shows the transient state in which the open-eye is still expanding. In the lower flow rate, we observe that the air bubbles rise in a cylindrical bubble column, with little mutual interaction between the bubbles, whereas at larger flow rates bubble-induced turbulence leads to a conical bubble plume. Moreover, at 5 litres per minute the oil layer atomizes and a large number of oil droplets and ligaments are found in the bulk of the water phase, an effect that was also observed in the corresponding experiment.
Fig. 4 shows the same snapshots in a view from above, with only the oil interface portrayed in red, so that we visualize the evolution of the open eye. To quantify it, we have defined a dimensionless open eye area as
@@
with $A_{\mathrm{ol}}^{0}$ being the initial area of the top of the oil layer. The evolution of $A^{*}$ in time is plotted in Fig. 5 for the 4 flow rates.
Figure 3: Side view of 3 snapshot for the lowest and highest flow rates to show the hydrodynamics of the Kim & Fruehan case. Left side: $Q=0.5$ l/min after 3, 7.5 and 12 seconds of physical time. Right side: $Q=5$ l/min after 1.5, 6 and 8.5 seconds of physical time.
t Refer to caption
Figure 4: Top view of 3 snapshots for the lowest and highest flow rates to show the hydrodynamics of the Kim & Fruehan case. Left side: $Q=0.5$ l/min after 3, 7.5 and 12 seconds of physical time. Right side: $Q=5$ l/min after 1.5, 6 and 8.5 seconds of physical time.
@@ 3.3 Mass transfer results
In the experimental campaigns, the flux is computed indirectly by measuring the concentration in the water phase over time and performing a linear regression using a lumped capacitance model given by the equation [8]
@@
The time evolution of the global Sherwood number given by Eq. 2 is plotted in Fig. 6 for the four mass transfer simulations of the Kim & Fruehan case.
@@
The Sherwood number is initially very large and decreases over the first 2.5 seconds of simulated time, corresponding to the transient growth of a chemical boundary layer in the water phase, after which it reaches a quasi-steady state. Larger fluctuations for the global Sherwood number are observed at higher flow rates due to turbulence and atomisation of the oil layer, causing significant variations in the effective transfer area and surface renewal time [38]. The time-average and standard deviation of the Sherwood number in the quasi-steady regime are measured.
@@
in an annulus surrounding the open eye, with the maximum value of approximately 60,000. The mass flux observed in the annulus corresponds to transfer in a shear layer in the water phase. The concentration boundary layer separates at the end of the annulus and low mass transfer rates are observed in the horizontal region of the oil-water interface. The transfer rate in this region is expected to be controlled by low-amplitude interfacial turbulence. Interestingly, in the high flow rate case, it is observed that the very small oil drops have a relatively small effect on the global Sherwood number, which could be due to a combination of saturation and that the density ratio between oil and water is small such that buoyancy would have little effect on convection. It is measured that the total contribution to mass transfer of the detached droplets was approximately 9 percent of the total mass transfer rate for a flow rate of 5 L/min.
Importantly, a Sherwood number of 50,000 corresponds to a concentration boundary layer of approximately 64 times smaller than the used grid size. This highlights the effectiveness of the SGS mass transfer model.
@@
To quantify the spatial distribution of species transfer between the bulk oil and water layers, we estimate a time-averaged local Sherwood number for each of the four simulations over a series of 20 snapshots, corresponding to 2 seconds of physical time. In Fig. 9 the projection on the $xz$ plane of this average Sherwood number $\overline{Sh}_{xz}$ is shown, with the colours indicating the average value of $\overline{Sh}_{xz}$ and the white zone depicting the average position of the open eye. It is observed that the value of the average Sherwood number in the annulus is weakly dependent on the flow rate and that the increase in the mass transfer rate as a function of flow rate is more of a consequence of the increase in effective transfer area due to the expanded open-eye.
@@ 3.4 Comparison of the results between simulations and experiments
In this section, we compare simulation results with analogous results measured by Joubert and by Kim & Fruehan in their experimental campaigns. This comparison is done through two macroscopic variables at pseudo-equilibrium: the time-averaged Sherwood number and open eye area. The results are plotted in Figs. 10 and 11.
We average the simulated open-eye area seen in Fig. 5 in the steady state regime. For the simulation results of Joubert first reported in [28], we plot the simulation results with the thin-film meniscus removed in post-processing, as the effective grid resolution without the meniscus is comparable to that used in the Kim & Fruehan simulations. The experimental open eye area was not reported directly in the Kim & Fruehan article but is calculated from the reported planar interfacial area.
The time-averaged Sherwood number is computed by averaging the the results for the global Sherwood number reported in Fig. 6 for each flow rate in the quasi-steady regime ($t>2.5$ s). The global Sherwood number is calculated for the Kim & Fruehan case using Eq. 2 and the reported values for $K_{w}A$.
@@
Figure 11: Simulated time-averaged global Sherwood number $\overline{Sh}$} against modified Froude number $N$ in compared with the experiments in [8, 6]. The standard error is used as error bars.
@@
The open-eye results portrayed in Fig. 10 demonstrate good agreement with both the Joubert and Kim & Fruehan experiments. As discussed previously, we confirmed that the simulations predicts the onset of atomisation of the oil layer. It is thus demonstrated that the simulation techniques used in this work accurately capture the global hydrodynamic behaviour in the metallurgical ladle. We must qualify our results for the open-eye area with the known problem of the capillary thin oil film spreading on water that reduces the open-eye size and constitutes a source of error on likely both the experimental and simulated open-eye measurements. This film is difficult to capture in the experiments, since it is not clearly distinguishable in the acquired pictures, while in the simulations it leads to mesh dependent results. In the current simulations, we remove the meniscus from our measured simulated open-eye with post-processing, but in the future it would be better to set a minimum thin-film thickness using the Manifold Death algorithm [39], which perforates thin sheets in an interface described by a VOF field at a user-defined thickness. The film-thickness parameter would be chosen from an estimated measurable film-thickness in experiments. According to the surface tension values for the argon-slag-steel system reported in [15, 5], slag is not expected to spread on liquid steel, thus removing this source of error when simulating the industrial system, but also showing a difference between the industrial and water experiments.
@@
From Fig. 11, we observe the same qualitative behaviour between the experiments with approximately the same slope of the dependence of $\overline{Sh}$ on $N$ in the first regime of mass transfer. It is observed in the Joubert case that the simulations predict a change in mass transfer regime near the experimental critical Froude number, though the mass flux in the second regime is under-predicted compared to experiments. Quantitative agreement is observed between the simulations and experiments in the first regime of mass transfer for the Joubert case, whereas the simulations over-predict the mass transfer rate at low flow rates compared to experiments for the Kim & Fruehan case. The over-prediction can be attributed to factors in both the simulations and the experiment. In the case of the former, the use of the error function as the boundary layer profile leads to a maximum in the predicted flux compared to the possible profiles from Eq. 18, with the other extreme Lévêque profile expected to decrease the predicted flux by approximately 12 percent [28]. It is also possible that the relatively coarse grid under-predicted the viscous dissipation in the oil phase. Lastly the difference in measured time between experiments (hours) versus simulations (seconds) means it is possible we are missing effects that occur at large time-scales e.g. the saturation of the oil layer. In the case of the experiments the inconsistent fluid volumes and missing measured surface tension and oil diffusivity lead to a likely small source of error. The most plausible explanation for the difference between the simulation and experiments is the position of the concentration probe used in the experiment, described in [6] to be 1 cm away from the bottom and near the wall of a model ladle. It seems likely that the tracer is not well-mixed in this region, with the concentration evolution of the probe lagging behind the average concentration evolution in the ladle, thus introducing an error when using the lumped capacitance model. In future experimental/simulation campaigns, the concentration in the water phase should be measured in experiments with passive measurement techniques (e.g. laser interferometry) in several places in the water such that an uncertainty on the experimental mass flux can be included. It is also possible, though potentially less preferable, to include simulated concentration probes directly in the CFD simulations.
4 Perspectives
The shown results demonstrate a significant advancement on the state of the art for computational mass transfer analysis in reduced-scale metallurgical ladles. It is demonstrated that combining multiscale approaches, with adaptive mesh resolution (AMR) to treat ranges of length scales in the hydrodynamic system combined with a subgrid-scale (SGS) model to treat the thin concentration boundary layers, relative agreement between numerics and simulations is achieved for the first time without extrapolating results simulated with artificially high diffusivities. The SGS model allows us to simulate very large $Sc$ numbers with the simulation cost comparable to simulating just hydrodynamics without mass transfer.
The simulations give much richer data than what is measurable in experiments. For example, we can directly measure the time-averaged velocity field to understand physical phenomena present in the system rather than rely on macroscopic, optically-measurable quantities such as the open-eye size. We can also directly measure the mass flux at the oil-water interface, which is potentially more reliable than the results produced by a concentration probe and the lumped-capacity model. We can use this rich data to develop a phenomenological model that predicts the measured quantities as a function of the flow rate. A mechanistic model that explains the measured quantities and attempts to unify all hydrodynamic and mass transfer experiments is under development and will be the subject of a future paper.

### v2

@@ Abstract
We use direct numerical simulation (DNS) to investigate mass transfer between liquid steel and slag during a metallurgical secondary refinement process through two reduced-scale water experiments, which reproduce the dynamics seen in an industrial bottom-blown ladle. A container is filled with water and topped by a thin layer of oil, representing the molten steel and slag, respectively. The system is agitated by a bubble plume that impinges on the oil layer and forms an open-eye. A tracer species, dissolved in the water, acts as a passive scalar that is progressively absorbed into the oil layer. Both the hydrodynamics and mass transfer in the system are studied and compared with experiments from the literature for ladles of different size and geometry.
The numerical simulation of mass transfer is challenging due to the high Péclet number, leading to extremely thin species boundary layers at the interface. Resolving the boundary layer is prohibitive even with adaptive grid techniques. A subgrid-scale (SGS) boundary layer model was employed to correct the scalar transport equation, allowing us to solve convection-dominated transport on relatively coarse grids. The hydrodynamics is investigated, and we analyze how the resultant flow field governs mass transport. The numerical results recover two flow regimes: a quasi-steady regime at low flow rates with small deformations of the oil-water interface and an atomising regime at large flow rates. Interfacial species transport is determined to be dominated in an annulus surrounding the open eye caused by a shear layer at the oil-water interface. It is observed that we achieve grid-independent macroscopic quantities that match relatively well with those observed in experiments, allowing use of simulation techniques as a complementary tool going forward.
@@ 1 Introduction
Performing experimental trials in the metallurgical industry is challenging because of high temperatures and the opaque nature of molten metals, consequently requiring specialized measurement techniques. For these reasons, industrial trials on the actual reactors are very expensive and practically unfeasible. Experimental campaigns are hence generally performed on reduced-scale model ladles and then a similarity criterion is used to grant a relation with the larger industrial ladles. While some studies use liquid metals at lower temperatures [3, 4, 5] some others further simplify the problem using fluids such as water and oil at room temperature [6, 7, 8]. In this case, a tracer species that is miscible in oil and water is chosen that mimics the species transfer kinetics observed in the industrial ladle. Thymol is used as the diffusive species in the experiments, as it miscible in both water and the chosen oil, has an equilibrium partition ratio between oil and water similar to that for sulfur between slag and steel, and interfacial mass transfer rate is rate-limited by water-side diffusion, such that the water system mimics the industrial system[6].
Measuring local quantities in the fluid system requires more sophisticated measurement techniques, so the flow has historically been studied by measuring and analyzing a few global quantities: the open-eye size, the fragmentation of the slag layer and the rate of mass transfer. The latter has been thoroughly investigated by Conejo [9] and Liu[10] by studying the effect of various physical and geometrical parameters on the mass transfer. Numerical simulations, on the other hand, can greatly improve the understanding of the fluid dynamics in gas-stirred ladles since they allow for easier measurements of local quantities and better visualisation of the flow behaviour. While the first simulations on the topic date more than 20 years back [11, 12], with the computational power available today increasingly complex simulations have become possible [13]. In particular, while the earliest simulations mainly focused on the flow patterns in the ladle, linking it to the mass transfer by means of models and theories, more recent studies started to address the mass transfer problem itself [14, 8]. There are however still some gaps in the understanding of the problem. Furthermore, most of the studies involving numerical simulations do not resolve the turbulent flow explicitly, but they rather use turbulence models such as large eddy simulations or Reynolds-averaged Navier-Stokes to treat smaller scale phenomena. Only recently some studies on ladles have been performed using direct numerical simulations (DNS) that solve turbulence to the smallest scales [15, 16]. Lastly, the majority of numerical studies has focused on the hydrodynamic problem and ignores resolving mass transfer.
@@ 2.1 Model description
$\overline{Sh}=\frac{K_{\mathrm{w}}Ah_{\mathrm{w}}}{L_{\mathrm{x}}^{2}D_{\mathrm{w}}},$ (2)
@@ 2.2 Mathematical model
The experiments analyzed are isothermal and they involve three immiscible phases: water, air and oil. While water and oil are incompressible, the injected air bubbles are in theory compressible. To simplify the model and reduce the computational load, we neglect the compressibility effects of air. This hypothesis is reasonable if the hydrostatic pressure of the system is significantly smaller than the atmospheric pressure $\rho_{\mathrm{w}}gh_{\mathrm{w}}\ll p_{\mathrm{atm}}$.
Esmaeeli & Tryggvason [20] and Cano-Lozano et al. [21] studied the effect of the numerical value of the air density in bubbly flows. They reported that a higher numerical air density had only a small influence on the rise velocity, which is explained by the pressure gradient in the bubble remaining much smaller than in the surrounding liquid. We also show in Appendix A that momentum-driven effects should have negligible effect on the vertical velocity of the bubble plume. Consequently, we can increase the air density with negligible effect on the flow. A reduction of the density ratio $\rho_{\mathrm{w}}/\rho_{\mathrm{a}}$, makes the pressure-Poisson problem converge significantly faster, increasing computational efficiency. We choose to use a value of the air density $\rho_{\mathrm{a}}$ that is ten times bigger than the air density at STP, as shown in Table 1.
Interface capturing is included with the multi-VOF approach, with three indicator functions $f_{i}$ to identify the different phases. The $f_{i}$ have to respect a volume conservation equation
@@
Under the formulated hypotheses, we write the incompressible Navier-Stokes equations for the three-phase problem using the one-fluid method as
@@
The spatial and temporal schemes used for the incompressible Navier-Stokes equations are described in [18, 19, 23]. It uses an operator split approach for the momentum equation, with a projection method used to enforce the continuity equation. The advection of each VOF scalar field is performed using the dimension-split geometric scheme of [25], which conserves mass to machine tolerance for a divergence-free velocity field.
@@
Furthermore, Henry’s law holds, stating that the ratio between the oil and water-side concentration is given by a partition ratio $P$
@@
Numerically, the scalar transport equation is solved using a two-field approach, where we transport the conserved quantities
@@
corresponding to the species mass in the oil and water phases. Two-scalar transport equations are solved, one corresponding to mass transfer in the water phase and the other in the oil phase. The phasic scalar transport equation for $g$ is written as
@@
where $A_{i}$ is the phasic area. The advection term is treated with a geometric scheme consistent with VOF advection and prevents numerical diffusion across the interface, as shown in reference [26]. The phasic diffusion term is treated with a time-implicit scheme, and the diffusive flux across the interface is included as equal and opposite explicit source terms in the two phasic diffusion solvers. Species mass is conserved up to the tolerance of the pressure-Poisson solver which enforces continuity.
@@ 2.3 Simulations setup
Furthermore, the Joubert simulations underestimated the spreading force at the three-phase contact line, which caused the open-eye area to be over-predicted by up to 30 percent compared to experiments. The improved surface tension calculation due to the hybrid curvature calculation described in 2.2 led to the formation of a thin oil meniscus that significantly reduced the size of the open eye, physically consistent with the positive spreading coefficient $S_{o}=\sigma_{aw}-\sigma_{ao}-\sigma_{ow}$ of the oil-water-air system. However, meniscus coverage was observed to be grid dependent, as coarse grids could not capture the thin film. Furthermore, comparisons to experiments would be difficult in the presence of thin film, as sophisticated optical tools were not used when measuring the open-eye in experiments. Using the meniscus area coverage as an error bar on the open-eye area in simulations led to relative agreement with experiments (see [34]).
@@ 2.3.4 Passive Scalar Subgrid-Scale modelling
To derive a subgrid model, we first assume that the mass transfer resistance in the water layer is much larger than that of the oil layer, known from the global kinetics of the system [6]. This allows us to compute the interfacial flux with a one-sided subgrid model in the water phase. We consider an advection-diffusion problem entirely located inside the layer $|z-z_{ow}|\ll\delta_{u,w}$, solved in the reference frame of the moving interface. In this sublayer, the velocity field $U$ may be linearized so that $U_{r}(z,t)=U_{\Sigma}(t)+\omega(t)(z-z_{ow})$, representing an interface slip velocity and a shear dependent term. We ignore tangential diffusion. The described boundary layer approximation leads to the simplified advection-diffusion equation for the concentration field $C(r,z,t)$
@@ 3 Results
This section presents the simulation results for both hydrodynamics and mass transfer. Detailed results are provided for the Kim & Fruehan case, while only principal results are highlighted for the Joubert case, for which additional plots and figures are available in the supplementary materials.
@@ 3.2 Hydrodynamics results
Fig. 3 shows the conical ladle of the Kim & Fruehan case from a side view for the highest and lowest air injection rates investigated, namely 0.5 and 5.0 l/min, at three subsequent time steps, to illustrate the evolution of the system in time up to the reaching of quasi-steady regime. The air interface is colored in green and the oil interface in red. The first snapshot of each simulation shows the transient state in which the open-eye is still expanding. In the lower flow rate, we observe that the air bubbles rise in a cylindrical bubble column, with little mutual interaction between them, whereas at larger flow rates bubble-induced turbulence leads to a conical bubble plume. Moreover, at 5 litres per minute the oil layer atomizes and a large number of oil droplets and ligaments are found in the bulk of the water phase, an effect that was also observed in the corresponding experiment.
Fig. 4 shows the same snapshots in a view from above, with only the oil interface portrayed in red, so that we can visualize the evolution of the open eye. To quantify its evolution, we have defined a dimensionless open eye area as
@@
with $A_{\mathrm{ol}}^{0}$ being the initial area of the top of the oil layer. The evolution of $A^{*}$ in time is plotted in Fig. 5 for the 4 flow rates. For all flowrates, we can distinguish a transient state in which the open eye area increases steadily and then the quasi-steady regime in which its value oscillates around a plateau value with the amplitude of the oscillations depending on the flow rate.
Figure 3: Side view of 3 snapshot for the lowest and highest flow rates to show the hydrodynamics of the Kim & Fruehan case. Top row: $Q=0.5$ l/min after 3, 7.5 and 12 seconds of physical time. Bottom row: $Q=5$ l/min after 1.5, 6 and 8.5 seconds of physical time.
H Refer to caption
Figure 4: Top view of 3 snapshots for the lowest and highest flow rates to show the hydrodynamics of the Kim & Fruehan case. Top row: $Q=0.5$ l/min after 3, 7.5 and 12 seconds of physical time. Bottom row: $Q=5$ l/min after 1.5, 6 and 8.5 seconds of physical time.
@@ 3.3 Mass transfer results
In the experimental campaigns, the flux is computed indirectly by measuring the concentration in the water phase over time and performing a linear regression using a lumped capacitance model [8], given by the equation
@@
The time evolution of the global Sherwood number given by Eq. 2 is plotted in Fig. 6 for the four mass transfer simulations of the Kim & Fruehan case. The Sherwood number is initially very large and decreases over the first 2.5 seconds of simulated time, corresponding to the transient growth of a chemical boundary layer in the water phase, after which it reaches a quasi-steady state. Larger fluctuations for the global Sherwood number are observed at higher flow rates due to turbulence and atomisation of the oil layer, causing significant variations in the effective transfer area and surface renewal time [44]. The time-average and standard deviation of the Sherwood number in the quasi-steady regime are measured.
@@
in an annulus surrounding the open eye, with the maximum value of approximately 60,000. The mass flux observed in the annulus corresponds to transfer in a shear layer in the water phase. The concentration boundary layer separates at the end of the annulus and low mass transfer rates are observed in the horizontal region of the oil-water interface. The transfer rate in this region is expected to be controlled by low-amplitude interfacial turbulence. Interestingly, in the high flow rate case, it is observed that the very small oil drops have a relatively small effect on the global Sherwood number, which could be due to a combination of saturation and that the density ratio between oil and water is small such that buoyancy would have little effect on convection. It is measured that the total contribution to mass transfer of the detached droplets was approximately 9% of the total mass transfer rate for a flow rate of 5 L/min. This was computed using the connected-components labelling algorithm described in [45, 46], used to segregate topologically distinct regions in a binary field. The bulk transfer, corresponding to the oil region with the largest volume, was thus discriminated from the detached droplets (all other regions). It is shown in B that the evolution of total oil-water interface area with flow rate does not correlate well with the observed mass transfer rates for the chosen flow rates. In a follow-up work, it could be interesting to investigate how the contribution to mass transfer of small droplets changes at higher flow rates.
It also important to point out that a Sherwood number of 50,000 corresponds to a concentration boundary layer that is approximately 64 times smaller than the used grid size. This highlights the effectiveness of the SGS mass transfer model.
@@
To quantify the spatial distribution of species transfer between the bulk oil and water layers, we estimate a time-averaged local Sherwood number for each of the four simulations over a series of 20 snapshots, corresponding to 2 seconds of physical time. In Fig. 9 the projection on the $xz$ plane of this average Sherwood number $\overline{Sh}_{xz}$ is shown, with the colors indicating the average value of $\overline{Sh}_{xz}$ and the white zone depicting the average position of the open eye. It is observed that the value of the average Sherwood number in the annulus is weakly dependent on the flow rate and that the increase in the mass transfer rate as a function of flow rate is more of a consequence of the increase in effective transfer area due to the expanded open-eye.
@@ 4.1 Comparison of the results between simulations and experiments
In this section, we compare simulation results with analogous results measured by Joubert and by Kim & Fruehan in their experimental campaigns. This comparison is done through two macroscopic variables at pseudo-equilibrium: the time-averaged Sherwood number and the open eye area. The results are plotted in Figs. 10 and 11.
We average the simulated open-eye area seen in Fig. 5 in the steady state regime. For the simulation results of Joubert, first reported in [34], we plot them with the thin-film meniscus removed in post-processing, as the effective grid resolution without the meniscus is comparable to that used in the Kim & Fruehan simulations. The experimental open eye area was not reported directly in the Kim & Fruehan article but is calculated from the reported planar interfacial area.
The time-averaged Sherwood number is computed by averaging the results for the global Sherwood number reported in Fig. 6 for each flow rate in the quasi-steady regime ($t>2.5$ s). The global Sherwood number is calculated for the Kim & Fruehan case using Eq. 2 and the reported values for $K_{w}A$.
@@
Figure 11: Simulated time-averaged global Sherwood number $\overline{Sh}$} against modified Froude number NN compared with the experiments in [8, 6]. The standard errors are used as error bars.
@@
The open-eye results portrayed in Fig. 10 demonstrate good agreement with both the Joubert and Kim & Fruehan experiments. As discussed previously, we confirmed that the simulations predicts the onset of atomisation of the oil layer. It is thus demonstrated that the simulation techniques used in this work accurately capture the global hydrodynamic behaviour in the metallurgical ladle. We must qualify our results for the open-eye area with the known problem of the capillary thin oil film spreading on water that reduces the open-eye size and constitutes a source of error on likely both the experimental and simulated open-eye measurements. This film is difficult to capture in the experiments, since it is not clearly distinguishable in the acquired pictures, while in the simulations it leads to mesh dependent results. In the current simulations, we remove the meniscus from our measured simulated open-eye with post-processing, but in the future it would be better to set a minimum thin-film thickness using the Manifold Death algorithm [47], which perforates thin sheets in an interface described by a VOF field at a user-defined thickness. The film-thickness parameter would be chosen from an estimated measurable film-thickness in experiments. We also have to point out that, according to the surface tension values for the argon-slag-steel system reported in [17, 5], slag is not expected to spread on liquid steel, thus removing this source of error when simulating the industrial system, but also showing a difference between the industrial and water experiments.
@@
From Fig. 11, we observe the same qualitative behaviour between the experiments and simulations with approximately the same slope of the dependence of $\overline{Sh}$ on $N$ in the first regime of mass transfer. It is observed in the Joubert case that the simulations predict a change in mass transfer regime near the experimental critical Froude number, though the mass flux in the second regime is under-predicted compared to the experiments. Quantitative agreement is observed between the simulations and experiments in the first regime of mass transfer for the Joubert case, whereas the simulations over-predict the mass transfer rate at low flow rates compared to experiments for the Kim & Fruehan case. The over-prediction can be attributed to factors in both the simulations and the experiment. In the case of the former, the use of the error function as the boundary layer profile leads to a maximum in the predicted flux compared to the possible profiles from Eq. 18, with the other extreme Lévêque profile expected to decrease the predicted flux by approximately 12 percent [34]. It is also possible that the relatively coarse grid under-predicted the viscous dissipation in the oil phase. Lastly the difference in measured time between experiments (hours) versus simulations (seconds) means it is possible we are missing effects that occur at large time-scales e.g. the saturation of the oil layer. In the case of the experiments the inconsistent fluid volumes and missing measured surface tension and oil diffusivity lead to a likely small source of error. The most plausible explanation for the difference between the simulation and experiments is the position of the concentration probe used in the experiment, described in [6] to be 1 cm away from the bottom and near the wall of a model ladle. It seems likely that the tracer is not well-mixed in this region, with the concentration evolution of the probe lagging behind the average concentration evolution in the ladle, thus introducing an error when using the lumped capacitance model. In future experimental/simulation campaigns, the concentration in the water phase should be measured in experiments with passive measurement techniques (e.g. laser interferometry) in several places in the water such that an uncertainty on the experimental mass flux can be included. It is also possible, though potentially less preferable, to include simulated concentration probes directly in the CFD simulations.
4.2 Scalability of the model